OpenStudy (anonymous):

subtract by writting the fractions with common denominaters.1/2-3/8

OpenStudy (anonymous):

Look for the lowest common multiple: 2: 2 4 6 [8] 10 12 ... 8: [8] 16 24 ... \[\frac{1}{2} = \frac{4}{8}\] So you can rewrite your original equation as \[\frac{4}{8} - \frac{3}{8}\] which equals \[\frac{1}{8}\]